OpenStudy (anonymous):

8m + 19 - m when m + 3. 8m + 19 - m when m + 3. @Mathematics

OpenStudy (anonymous):

you mean m = 3

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

40

OpenStudy (diyadiya):

8(3)+19-(3) =24+19-3 =43-3 =40

OpenStudy (anonymous):

40